### v2

# GUARD: Glocal Uncertainty-Aware Robust Decoding for Effective and Efficient Open-Ended Text Generation

Abstract
Open-ended text generation faces a critical challenge: balancing coherence with diversity in LLM outputs. While contrastive search-based decoding strategies have emerged to address this trade-off, their practical utility is often limited by hyperparameter dependence and high computational costs. We introduce GUARD, a self-adaptive decoding method that effectively balances these competing objectives through a novel "Glocal" uncertainty-driven framework. GUARD combines global entropy estimates with local entropy deviations to integrate both long-term and short-term uncertainty signals. We demonstrate that our proposed global entropy formulation effectively mitigates abrupt variations in uncertainty, such as sudden overconfidence or high entropy spikes, and provides theoretical guarantees of unbiasedness and consistency. To reduce computational overhead, we incorporate a simple yet effective token-count-based penalty into GUARD. Experimental results demonstrate that GUARD achieves a good balance between text diversity and coherence, while exhibiting substantial improvements in generation speed. In a more nuanced comparison study across different dimensions of text quality, both human and LLM evaluators validated its remarkable performance. Our code is available at https://github.com/YecanLee/GUARD.
GUARD: Glocal Uncertainty-Aware Robust Decoding for Effective
and Efficient Open-Ended Text Generation
Yuanhao Dingthanks:   Equal contribution1    Esteban Garces Arias11footnotemark: 12,3    Meimingwei Li11footnotemark: 12    Julian Rodemann2,4 Matthias Aßenmacher2,3    Danlu Chen5    Gaojuan Fan1    Christian Heumann2 Chongsheng Zhangthanks:   Corresponding author1 1Henan University, 2Department of Statistics, LMU Munich 3Munich Center for Machine Learning (MCML) 4CISPA Helmholtz Center for Information Security, Saarbrücken 5University of California, San Diego {yhding, fangaojuan, cszhang}@henu.edu.cn {esteban.garcesarias, matthias}@stat.uni-muenchen.de, M.Li@campus.lmu.de {J.Rodemann, Christian.Heumann}@lmu.de, danlu@ucsd.edu
1 Introduction
Neural text generation models based on the Transformer decoder (Vaswani et al., 2017) have revolutionized natural language generation tasks such as story creation (Fang et al., 2021), context completion (Leng et al., 2024), or dialogue generation (Su et al., 2021). Decoding strategies play a crucial role in text generation with large language models (LLMs), affecting both the quality of the generated text and the computational efficiency during inference. The most commonly used strategies can be broadly categorized in deterministic Freitag and Al-Onaizan (2017); Carlsson et al. (2025) and stochastic approaches Fan et al. (2018); Nguyen et al. (2025a); Aichberger et al. (2025). Deterministic approaches are likelihood-based techniques that typically (over-)emphasize coherence at the cost of diversity and are prone to producing repetitive text, a phenomenon referred to as degeneration. Stochastic approaches, on the other hand, aim to improve the diversity of texts, potentially sacrificing coherence and leading to semantic inconsistencies (Welleck et al., 2020).
This inherent coherence vs. diversity trade-off has motivated recent work to develop decoding strategies able to balance these two competing objectives. In particular, Contrastive Search (CS; Su et al., 2022) introduced a weighted combination of model confidence (favoring coherence) and a degeneration penalty (promoting diversity) controlled by fixed, tunable hyperparameters throughout the generation. Adaptive Contrastive Search (ACS; Garces-Arias et al., 2024) addressed this hyperparameter dependence by dynamically adjusting them based on the local uncertainty of the model. In doing so, this approach still has its limitations as it relies only on local uncertainty information, and it substantially increases the computational overhead at inference time. In this study, we aim to address the following research questions (RQ):
  1. 1.
    RQ1: Can we improve text quality in open-ended text generation compared to existing approaches (§4.2)?
  2. 2.
    RQ2: Can we smooth out abrupt entropy deviations during generation without compromising statistical properties such as unbiasedness and consistency (§3.1)?
  3. 3.
    RQ3: Can we substantially reduce the latency and computational cost compared to (A)CS (§4.2, "Generation Speed")?
To address the above questions, we introduce a novel concept called "Glocal" uncertainty, which combines global entropy estimates with local entropy deviations to integrate long-term and short-term uncertainty signals, thereby maintaining robust decoding while ensuring statistical unbiasedness and consistency. Based on Glocal uncertainty, we propose a novel decoding strategy, Glocal Uncertainty-Aware Robust Decoding (GUARD), which effectively balances the coherence and diversity of the generated texts (RQ1), addresses the smoothness of entropy estimates (RQ2), and improves inference speed (RQ3).
Our Contributions can be summarized as follows:
  1. 1.
    We introduce "Glocal" uncertainty, which integrates global entropy and instantaneous uncertainty signals to smooth out strong deviations in entropy, resulting in more robust decoding. We provide statistical guarantees (unbiasedness and consistency) for the proposed global entropy estimator, which is an integral part of the Glocal uncertainty.
  2. 2.
    Based on Glocal uncertainty, we design GUARD, a self-adaptive decoding strategy that effectively and efficiently balances coherence and diversity. Compared to (A)CS, it reduces the computational cost remarkably by incorporating a simple, yet effective, token count penalty.
  3. 3.
    We evaluate GUARD through comprehensive experimentation that spans diverse models, datasets, and decoding configurations, comparing it against state-of-the-art (SOTA) decoding strategies. Our assessment employs automatic, human, and LLM evaluations, which consistently show that GUARD represents a high-performing decoding strategy.
2 Related work
Deterministic Strategies
primarily rely on maximizing the sequence probability. Greedy Search selects tokens based on pure maximum likelihood, while Beam search Freitag and Al-Onaizan (2017) improves upon this by maintaining multiple candidates simultaneously, ultimately yielding the most probable sequence. Further research (Vijayakumar et al., 2018; Holtzman et al., 2020; Garces Arias et al., 2025) has demonstrated that this often results in text degeneration and repetitive patterns. Recent developments, such as hyperfitting Carlsson et al. (2025), show promising solutions to this problem, even within greedy search frameworks.
Figure 1: GUARD leverages local and global Shannon entropy deviations as proxies for model uncertainty, automatically adjusting CS parameters and incorporating a token count penalty for next-token selection.
Stochastic Strategies
based on sampling have emerged as a response to the limitations of deterministic approaches, particularly addressing text monotony and degeneration. Temperature sampling Ackley et al. (1985) introduces a hyperparameter $\tau$ that modulates the sharpness of the output distribution, enabling control over generation diversity. Top-$k$ (Fan et al., 2018) and Top-$p$ sampling Holtzman et al. (2020) truncate the output distribution to avoid sampling from less reliable tail probabilities. Typical sampling Meister et al. (2023a) constrains the sampling distribution to tokens whose negative log-probabilities fall within a specific range of the model’s conditional entropy, while min-$p$ sampling Nguyen et al. (2025a) introduces dynamic truncation based on the model’s confidence. Adaptive Decoding (AD, Zhu et al., 2024) employs an entropy-based confidence metric to optimize candidate selection. While these sampling methods effectively reduce degeneration, they risk inconsistency, compromising overall text quality Basu et al. (2021).
Contrastive Strategies
typically aim at balancing coherence and diversity in text generation tasks. Contrastive Decoding (CD, Li et al., 2023) leverages the differential between expert and amateur language models to select tokens that maximize their log-likelihood difference. CS evaluates top-$k$ tokens using a combination of model confidence and degeneration penalty to enhance output diversity. ACS builds upon CS by eliminating hyperparameters, addressing the unexplored impact of hyperparameter selection on generation quality Garces Arias et al. (2025), at the cost of additional computational overhead. Moreover, its focus on immediate uncertainty may overlook temporal fluctuations that could benefit from more robust generation strategies Chen et al. (2025). These limitations motivate GUARD, a novel decoding strategy that offers an efficient, self-adaptive decoding guided by uncertainty.
3 Methodology
Figure 1 illustrates GUARD, which (a) leverages global entropy $H_{\text{glob, $t$}}$ (Eq. (1)) to capture model uncertainty fluctuations throughout the decoding process and (b) introduces Glocal uncertainty (denominator in Eq. (6)). It thus considers both long- and short-term uncertainty fluctuations to enable adaptive control over $k_{t}$ and $\alpha_{t}$ during (A)CS for the top-$k_{t}$ most semantically accurate tokens. The method incorporates a simple yet effective token count-based penalty for repeated tokens.
Prompt: A shiver ran through her and she walked back inside. There was a chill in the air and she was only wearing a t-shirt and jeans. The
Generated story: house was warm, though. Her mother had probably turned on another heater while they were out.
Figure 2: Left: Local (solid) and global (dashed/interval) Shannon entropy over the time steps of the generation. Right: Strong changes are smoothed to provide a robust entropy estimation before computing $k$ and $\alpha$ over time, leading to increased stability.
3.1 Global Entropy and Statistical Properties
In the related ACS strategy, the Shannon entropy $H(X)_{t}=-\sum_{x\in\mathcal{V}}p(x\mid\boldsymbol{x}_{<t})\ln p(x\mid\boldsymbol{x}_{<t})$ is used to dynamically adjust hyperparameters at the time step $t$. As is customary, $X$ denotes a discrete random variable with $\mathcal{V}$ referring to the tokens; $t\in\{1,\dots,T\}$ denotes a single generation step, limited to the maximum sequence length (e.g., $T=256$).
[…]
$H_{\text{glob},t}=\frac{\sum_{i=1}^{t}(\lambda^{t-i}\cdot H(X)_{i})}{\sum_{i=1}^{t}\lambda^{t-i}}$ (1)
The parameter $\lambda\in(0,1]$ denotes a weighting coefficient. We employ $\lambda^{t-i}$ to reflect the temporal evolution of model uncertainty over time (i.e., over sequential generation steps), which can dynamically adjust the contribution of each generation step to $H_{\text{glob},t}$.
In the following, we will provide Proposition 1 to prove that $H_{\text{glob},t}$ is an unbiased estimator of the instantaneous entropy $H(X)_{t}$, irrespective of the concrete choice of $\lambda\in(0,1]$. In Proposition 2, we further prove that $H_{\text{glob},t}$ is a consistent estimator of $H(X)_{t}$ under reasonable assumptions. This tells us that $H_{\text{glob},t}$ converges in probability, i.e., as $t\rightarrow\infty$, to the instantaneous entropy. Thus, it is not only unbiased but also has diminishing variance when the generation is sufficiently long. Detailed proofs of both results can be found in Appendix A.
Proposition 1 (Unbiasedness).
Assume that the process $\{H(X)_{t}\}_{t\geq 1}$ is stationary. Then,
$\mathbb{E}\bigl{[}H_{\text{glob},t}\bigr{]}=\mathbb{E}[H(X)_{t}]$ (2)
for any $\lambda\in[0,1)$.
Sketch of Proof.
The result directly follows from the expectation’s linearity and $\{H(X)_{t}\}_{t\geq 1}$ being stationary. Linearity gives
$\mathbb{E}\bigl{[}H_{\text{glob},t}\bigr{]}=\sum_{i=1}^{t}\frac{\lambda^{t-i}}{\sum_{s=1}^{t}\lambda^{t-s}}\,\mathbb{E}\bigl{[}H(X)_{i}\bigr{]}$ (3)
Stationarity implies
$\forall t\in\{1,\dots,T\}:\;\mathbb{E}\bigl{[}H(X)_{t}\bigr{]}=H_{0},$ (4)
[…]
Proposition 2 (Consistency).
Assume the process $\{H(X)_{t}\}_{t\geq 1}$ is stationary, ergodic and bounded. Further assume that the covariances $\operatorname{Cov}\bigl{(}H(X)_{t},H(X)_{s}\bigr{)}$ decay sufficiently fast as $|t-s|\to\infty$. Then, as $t\rightarrow\infty$,
$H_{\text{glob}}\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\mathbb{E}[H(X)_{t}]$ (5)
Sketch of Proof.
We have that all $H(X)_{t}$ are bounded and their covariances decay sufficiently fast as |ts||t-s| increases. Moreover, the weights form a convex combination. Under standard mixing conditions, we show that
$\operatorname{Var}\bigl{(}H_{\text{glob}}\bigr{)}=\mathcal{O}\Bigl{(}\frac{1}{t}\Bigr{)},$
[…]
$\mathbb{P}\Bigl{(}\bigl{|}H_{\text{glob}}-\mathbb{E}[H_{\text{glob}}]\bigr{|}>\epsilon\Bigr{)}\leq\frac{\operatorname{Var}\bigl{(}H_{\text{glob}}\bigr{)}}{\epsilon^{2}}\to 0$
as $t\to\infty$, from which the claim follows. $\square$.
Brief Summary.
In Proposition 1 and Proposition 2, we prove that $H_{\text{glob},t}$ is not only unbiased, but also has diminishing variance when the generation is long enough. Such statistical properties support that we can reasonably estimate the instantaneous entropy with $H_{\text{glob},t}$. This answers RQ2 – the intended smoothing of the uncertainty estimation does not come at the price of losing statistical validity. However, note that both results concern our internal estimators of the entropy, not the inferential properties of the final language model.
Interpretation.
Our method employs $\lambda^{(t-i)}$ to reflect the temporal evolution of model uncertainty. Specifically, as the time step $i$ approaches the current moment $t$, its contribution to $H_{\text{glob}}$ increases. Conversely, time steps distant from $t$ have diminishing effects on $H_{\text{glob}}$. The $\lambda^{(t-i)}$ factor implements a temporal decay mechanism, dynamically adjusting each time step’s contribution to $H_{\text{glob}}$. We later set $\lambda=0.95$, but recall that our method’s unbiasedness and consistency hold for any $\lambda\in(0,1]$. [^1] We experimentally verified that performance is relatively insensitive to the chosen value of $\lambda$ (cf. Tab. 6, Appendix B) and thus set $\lambda=0.95$ as an intermediate value. Therefore, our Global entropy measure can effectively estimate local uncertainties in a statistically unbiased and consistent way.
3.2 Glocal Uncertainty and Candidate Token Set
To determine $k_{t}$, that is, the selection of candidate tokens in CS, we introduce Glocal uncertainty. It simultaneously considers both the global and local uncertainty fluctuations in the token generation steps, denoted as $\delta_{glob}$ and $\delta_{loc}$, respectively. Specifically, we present Eq. (6) to automatically derive $k_{t}$ for the Top-$k_{t}$ candidate tokens.
$\scriptstyle k_{t}=10\cdot(1-\frac{1}{\exp\left(\lambda_{k}\cdot\delta_{\text{loc}}+(1-\lambda_{k})\cdot\delta_{\text{glob}}\right)+1})+5$ (6)
$\scriptstyle\delta_{\text{loc}}=q\cdot\operatorname{arctanh}\left(\frac{H(X)_{t}-\operatorname{med}\left(H(X)_{t-w:t-1}\right)}{\ln|\mathcal{V}|}\right)$ (7)
$\scriptstyle\delta_{\text{glob}}=q\cdot\operatorname{arctanh}\left(\frac{\operatorname{med}\left(H(X)_{t-w:t-1}\right)-\operatorname{med}\left(H_{\text{glob},t}\right)}{\ln|\mathcal{V}|}\right)$ (8)
$\lambda_{k}=\frac{|\delta_{\text{loc}}|}{|\delta_{\text{loc}}|+|\delta_{\text{glob}}|+\epsilon}$ (9)
where the constants 10 and 5 for the parameter $k_{t}$ define an effective range for exploration (Garces-Arias et al., 2024). $\delta_{\text{loc}}$ measures the model’s instantaneous uncertainty fluctuation compared to recent steps, i.e., the instantaneous uncertainty at the current $t^{th}$ step $H(X)_{t}$ versus the median uncertainty in the recent time window ww, [^2] To select the optimal value for $w$, we conducted extensive experiments, which are presented in Appendix B (Tables 814). We observe that $w$ does not notably influence text quality. Ultimately, we set $w=7$, keeping it constant for our experiments. denoted as $\operatorname{\textit{med}}\left(H(X)_{t-w:t-1}\right)$. $\delta_{\text{glob}}$ assesses the overall variation between the median of the short-term entropy $\operatorname{\textit{med}}\left(H(X)_{t-w:t-1}\right)$ and the median of the long-term uncertainty trend up to the current $t^{th}$ step $\operatorname{\textit{med}}\left(H_{\text{glob},t}\right)$. [^3] The choice of median and arctanh was experimentally validated. Our experiments compared different pairings, including winsorized mean, mean, median, logarithmic map, and arctanh, with detailed results shown in Table 7, Appendix B. $\lambda_{k}$ refers to the dynamic adjustment of local and global entropy during the generation process. $q$ is an adaptive temperature parameter influencing the dynamic adjustment magnitude of $k_{t}$ and $\alpha_{t}$. It adaptively adjusts by integrating changes in both current and global uncertainties:
$q=1.0+r_{\text{change}}+r_{\text{difference}}$ (10)
$r_{\text{change}}$ measures the rate of uncertainty change at $t$:
$r_{\text{change}}=\frac{|H(X)_{t}-H(X)_{t-1}|}{H(X)_{t-1}+\epsilon}$ (11)
$r_{\text{difference}}$ assesses the difference between local uncertainty within the current time window and the overall trend:
$r_{\text{difference}}=\frac{|\text{med}(H(X)_{t-w:t-1})-\text{med}(H_{\text{\text{glob},t}})|}{\text{med}(H(X)_{t-w:t-1})+\epsilon}$ (12)
When $r_{\text{change}}$ or $r_{\text{difference}}$ is large, $q$ increases, causing $\delta_{\text{loc}}$ and $\delta_{\text{glob}}$ to be amplified. $\lambda_{k}$ and $\lambda_{\alpha}$ dynamically allocate weights based on their ratio to adjust $k_{t}$ and $\alpha_{t}$, which are part of the token count penalty. When uncertainty is high, $k_{t}$ increases, potentially introducing repetitive candidate words, but also $\alpha_{t}$ increases, amplifying the degradation penalty and counteracting the effect of a higher $k_{t}$. Conversely, a decrease of $k_{t}$ and $\alpha_{t}$ results in outputs that favor high-confidence candidate words.
Rationale of Glocal Uncertainty.
In Eq. (6), the denominator $C_{\text{sum}}=\exp(\lambda_{k}\cdot\delta_{\text{loc}}+(1-\lambda_{k})\cdot\delta_{\text{glob}})$ can be regarded as an alternative to perplexity Jelinek et al. (1977), since the latter is the exponential of the total uncertainty in the prediction up to the current $t^{th}$ step, where Glocal uncertainty calculates the variations between the instantaneous entropy at the $t^{th}$ step with respect to short-term median entropy, as well as the short- vs. longer-term entropy variations. Similar to perplexity, a lower $C_{\text{sum}}$ value is preferred. When pronounced uncertainty variations occur, higher perplexity is indicated, suggesting we should select a larger $k_{t}$ value to expand the candidate token set, accommodating model instability. Conversely, when variations are minimal, a smaller $k_{t}$ value suffices, as the model demonstrates stability, allowing us to evaluate fewer candidate tokens.
Once $k_{t}$ is determined, we select the adequate Top-$k_{t}$ candidate tokens $\mathcal{V}^{(k_{t})}$, then design a token count penalty strategy to increase the diversity of the generated texts by reducing repetitiveness. Rather than employing a cosine-similarity-based penalty as in (A)CS, GUARD (as defined in Eq. (13)) circumvents the over-penalization of semantic similarity, leads to a more effective reduction of repetitions (and hence increased diversity), and accelerates the generation speed:
$\scriptstyle P(x_{next})=\underset{v\in\mathcal{V}^{(k_{t})}}{\arg\max}\Bigg{\{}{p_{\theta}(v\mid\boldsymbol{x}_{<k_{t}})}\times\underbrace{\alpha_{t}^{token\ counts(v)}}_{\text{degeneration penalty}}\Bigg{\}}$ (13)
where $\alpha_{t}$ is the penalty coefficient and $\text{token\ counts}(v)$ denotes the frequency of the candidate $v$ in the previously generated sequence. To dynamically determine $\alpha_{t}$, we reuse the core of Eq. (6), replacing $|\mathcal{V}|$ with $k_{t}$ in Eq. (7) and (8):
$\alpha_{t}=1-\frac{1}{\exp\left(\lambda_{k}\cdot\delta_{\text{loc}}+(1-\lambda_{k})\cdot\delta_{\text{glob}}\right)+1}$ (14)
When $\mathcal{V}^{(k_{t})}$ is small, each repetitive token in $\mathcal{V}^{(k_{t})}$ will receive a very light penalty, and vice versa. Appendix C illustrates the entire algorithm as pseudo-code to facilitate understanding of its flow. In Figure 2, we demonstrate that fluctuations can be smoothed to ensure a robust entropy estimation (left) when using our technique in determining $k_{t}$ and $\alpha_{t}$ over time (right). Our experiments show that the token count-based penalty strategy can substantially reduce repetitions, therefore increasing the diversity of the text outputs while enhancing text generation speed compared to using cosine similarity.
Self-Adaptive Design of GUARD.
GUARD achieves self-adaptivity for all quality-determining hyperparameters in CS methods; while it exhibits hyperparameters $\lambda$ and $w$, extensive testing (Appendix B) confirms that specific choices have negligible impact on generation quality, unlike other decoding methods (Garces Arias et al., 2025).
4 Experiments
4.1 Experimental Setup
This section will introduce the metrics and datasets used in our experiments, as well as the baseline decoding strategies we use for reference and the LLMs we employ.
Strategy Wikitext (n = 1314) Wikinews (n = 2000) BookCorpus (n = 1947)
Div. MAUVE Coh. BERT-Sc. Div. MAUVE Coh. BERT-Sc. Div. MAUVE Coh. BERT-Sc.
Greedy $16.97$ $48.28$ 0.78-0.78 $81.45$ $28.13$ $74.91$ 0.90-0.90 $82.51$ $4.07$ $17.99$ 0.65-0.65 $79.92$
Beam ($5$) $14.42$ $26.39$ 0.77-0.77 $81.20$ $24.94$ $54.31$ 0.87-0.87 $83.23$ $5.39$ $18.69$ 0.56-0.56 $80.03$
Temp. ($0.9$) $84.81$ $92.85$ -2.43 $81.86$ $92.71$ 96.29 2.41-2.41 $83.16$ $89.58$ $85.54$ 2.33-2.33 $81.78$
Top-$k$ ($50$) $82.69$ $88.97$ 1.98-1.98 $80.64$ $92.13$ $94.52$ 2.01-2.01 $83.17$ $91.54$ $93.24$ 2.52-2.52 $81.29$
Top-$p$ ($0.95$) $82.47$ $86.37$ 2.11-2.11 $81.50$ $91.47$ $96.09$ 2.15-2.15 $82.84$ $88.92$ 94.81 2.29-2.29 $81.31$
Typical $69.54$ $81.57$ 1.78-1.78 $81.34$ $82.31$ $95.66$ 1.83-1.83 83.86 $96.29$ $88.58$ 3.68-3.68 $80.06$
CD $72.70$ $79.76$ 2.61-2.61 $81.72$ $75.28$ $76.20$ 2.36-2.36 $82.57$ $76.00$ $72.10$ -2.53 $81.51$
AD $87.19$ 92.20 2.51-2.51 $81.83$ $91.49$ $94.38$ 2.35-2.35 $82.52$ $90.70$ $94.49$ 2.49-2.49 $81.82$
CS ($5,0.6$) $72.08$ $80.52$ 1.43-1.43 $81.53$ $87.47$ $91.47$ 1.31-1.31 $83.01$ $68.26$ $87.92$ 1.24-1.24 $81.20$
CS ($10,0.6$) $78.68$ $73.50$ 1.69-1.69 $80.57$ $91.38$ $90.49$ 1.57-1.57 $83.09$ $82.60$ $75.53$ 1.57-1.57 $81.22$
ACS ($q=1$) $81.57$ $78.82$ 1.95-1.95 $80.00$ 92.79 $91.23$ 1.72-1.72 $81.50$ $87.86$ $78.32$ 1.52-1.52 $79.79$
GUARD 92.86 $90.82$ 2.61-2.61 81.90 $95.20$ $93.60$ -2.38 $83.16$ 96.18 $92.59$ -2.53 81.83
Human $93.84$ $100.00$ 2.43-2.43 $100.00$ $93.50$ $100.00$ 2.86-2.86 $100.00$ $94.98$ $100.00$ 2.99-2.99 $100.00$
Table 1: Averaged automatic evaluation results for Qwen2.5-7B. Hyperparameters for competing strategies chosen based on experimental evaluation (Appendix G, Tables 18, 19, 20) or based on the original paper (CS, ACS). Scores closest to human highlighted in bold.
Dataset Human Evaluation
(n = 200 Sem. Coh.* (%)↑ Fluency** (%)↑
each) ACS GUARD ACS GUARD
Wikitext 37.50 62.50 43.75 56.25
Wikinews 40.00 60.00 45.25 54.75
BookCorpus 48.75 51.25 52.50 47.50
All 42.00 58.00 47.00 53.00
Table 2: Human evaluation: Share of human evaluators favoring a strategy w.r.t. perceived semantic coherence and fluency. P-values for the exact binomial test: $p_{\text{Sem. Coh.}}=2.875e-06$ , $p_{\text{Fluency}}=0.01819$. Best results are highlighted in bold.
Metric GUARD wins ACS wins Tie GUARD wins Top-$k$ wins Tie GUARD wins Top-$p$ wins Tie
Overall 34.3 30.6 35.1 65.8 17.5 16.7 49.2 34.2 16.7
Fluency 40.7 35.3 24.0 76.7 23.3 0.0 56.7 43.3 0.0
Sem. Coh. 40.7 34.0 25.3 80.0 20.0 0.0 58.3 41.7 0.0
Factuality 4.0 10.0 86.0 0.0 0.0 100.0 0.0 0.0 100.0
Informativeness 40.0 35.3 24.7 78.3 21.7 0.0 60.0 40.0 0.0
Interestingness 40.0 34.7 25.3 80.0 20.0 0.0 60.0 40.0 0.0
Story Development 40.7 34.0 25.3 80.0 20.0 0.0 60.0 40.0 0.0
Table 3: Win shares (%) of GUARD vs. ACS/Top-kk/Top-$p$ for LLM evaluation for $n=150$ continuations. Best results are highlighted in bold.
LLMs.
We employ six LLMs: Qwen2.5-7B Qwen (2024), Deepseek-llm-7B-base DeepSeek-AI (2024), Mistral-v0.3 Jiang et al. (2024), Llama-3 Dubey et al. (2024), Llama-2 at al. (2023), Gemma-7B Mesnard et al. (2024).
Evaluation Metrics.
We employ the following five commonly used metrics to automatically assess generation quality: Diversity Su et al. (2022), MAUVE Pillutla at al. (2021), Coherence Su et al. (2022), BERTScore Zhang et al. (2020). Formulas are provided in Appendix D.
Human Evaluation.
To evaluate the quality of the generated text, human evaluators assessed two key aspects: semantic coherence and fluency. Four native English speakers evaluated 600 pairs of competing text continuations, which were uniformly distributed across the three datasets. We measured inter-rater agreement using Fleiss’ Kappa and performed exact binomial tests to determine statistically significant differences between ACS and GUARD. Further details on the evaluation are provided in Appendix E.
LLM Evaluation.
We employ LLM-as-a-judge (judge: GPT-4) to systematically evaluate the quality of the generated text on $n=150$ continuations, uniformly distributed across the three datasets. This method aims to approximate the evaluation process of human experts on a large scale. For this, we measured six core metrics: fluency, coherence, factuality, informativeness, interestingness, and story development. The prompt is provided in Appendix F.
Datasets and Baselines.
We focus on three datasets for open-ended text generation from different domains: Wikinews (n = 2000), Wikitext (n = 1314; Merity et al., 2017), and BookCorpus (n = 1947; Zhu at al., 2015). We compare GUARD to SOTA LLM decoding methods from three categories: Deterministic (greedy search, beam search with $B\in\{3,5,10,15,20,50\}$), stochastic (temperature sampling with $\tau=0.9$, top-$k$ sampling with $k\in\{3,5,10,15,20,50\}$, top-$p$ sampling with $p\in\{0.60,0.70,0.80,0.90,0.95\}$, typical sampling with $\tau=0.2$, adaptive decoding), and contrastive approaches (CD, CS with $k\in\{5,10\}$, $\alpha=0.6$, and ACS with $q=1$). Across all experiments, we used a maximum token length of 256 tokens for the text generations.
Hardware.
We used NVIDIA 4090 (24GB memory) and NVIDIA H100 (80GB memory).
4.2 Results
Automatic evaluation results.
The automatic evaluations of different methods, based on diversity, MAUVE, and coherence, are presented in Table 1. We observe that, in general, GUARD outperforms all other competing strategies in both diversity and coherence. In terms of MAUVE, it is roughly on par with the best competing methods on each dataset. More importantly, it achieves results closest to human references without relying on specific hyperparameter choices. We observe that GUARD generally performs well according to BERTScore, which, however, is not very discriminative. Furthermore, we further validate the robustness of our method for other model architectures (cf. Appendix G, Tables 1617). Notably, GUARD is well-suited for decreasing the repetitive use of single tokens, as indicated by the high diversity values across all three benchmark datasets. While one might suspect that this improvement comes at the cost of coherence, our experiments show that GUARD also consistently exhibits coherence values that are very close to those of the human-written completions.
Metric Dataset Human Evaluation LLM Evaluation Agree?
($n=50$ each) GUARD (%) ACS (%) Ties (%) GUARD (%) ACS (%) Ties (%)
Sem. Coh. Wikitext 44.0 19.0 37.0 42.0 32.0 26.0 $\checkmark$
Wikinews 43.5 23.5 33.0 40.0 34.0 26.0 $\checkmark$
BookCorpus 38.0 35.5 26.5 40.0 36.0 24.0 $\checkmark$
Overall ($n=150$) 41.8 26.0 32.2 40.7 34.0 25.3 $\checkmark$
Fluency Wikitext 23.0 10.5 66.5 42.0 32.0 26.0 $\checkmark$
Wikinews 21.5 12.0 66.5 40.0 40.0 20.0
BookCorpus 13.0 18.0 69.0 40.0 34.0 26.0
Overall ($n=150$) 19.2 13.5 67.3 40.7 35.3 24.0 $\checkmark$
Table 4: Win shares (Human and LLM evaluation) w.r.t. coherence and fluency of GUARD and ACS. Best results are highlighted in bold.
Human evaluation.
We only compare ACS and GUARD for the sake of practicability and since both show strong performance (cf. Table 1. The results of this study are summarized in Table 2: Overall, GUARD demonstrates superior performance compared to ACS in terms of both semantic coherence and fluency, with the sole exception of fluency on the BookCorpus dataset. Specifically, human preferences indicate that GUARD outperforms ACS on average by approximately 6% in semantic coherence and 16% in fluency across all datasets. Both differences are statistically significant at the $\alpha=0.05$ level. The inter-rater reliability (Fleiss’ $\kappa$ = 0.41) shows moderate agreement among evaluators. In Appendix H, we include a small case study for the interested reader that further substantiates these findings. Overall, the human evaluation comprehensively addresses RQ1, confirming that both automatic metrics and human judgments favor our approach over alternative methods.
LLM evaluation.
We include LLM evaluation as an approximation of human judgments, comparing GUARD to its natural baseline ACS, as well as Top-$k$ and Top-$p$. The evaluation focuses on the six dimensions described in the experimental setup. The results (cf. Table 3) demonstrate that GUARD outperforms all three competitors in five out of six metrics (except factuality), where ACS has a slight edge, and comparison to the other two strategies results in a tie. When comparing human to LLM judgments for coherence and fluency (cf. Table 4), results (a) demonstrate strong agreement and (b) show that human evaluators indicate an even stronger preference for GUARD’s coherence compared to LLM-judges.
Method sec/story #tokens/sec #tokens/story
CS (10, 0.6) 11.6 22.0 249.7
ACS ($q$ = 1) 15.7 16.3 255.7
ACS ($q$ = 2) 15.9 16.1 255.8
ACS ($q$ = 8) 16.3 15.3 255.8
GUARD 4.42 28.3 255.9
Table 5: Average generation speed ($n=30$) of GUARD with CS and ACS ($q\in\{1,2,8\}$). Best results in bold.
Generation speed.
In Table 5, we compare GUARD to CS and ACS in terms of the average generation speed. Based on 30 generation samples, all three methods produce generations of comparable lengths. CS spends 11.6 seconds per story, ACS takes an even longer, from 15.7 seconds to 16.3 seconds (across different temperature settings), while GUARD only needs 4.42 seconds per story. We measure the average number of tokens decoded per second, and observe that CS decodes on average 21.98 tokens/second, ACS between 15.35 and 16.29, and GUARD decodes 28.3 tokens/second on average. This answers RQ3, since the generation speed of CS-based approaches can be substantially increased using GUARD.
5 Discussion
Stationarity Assumption.
[…]
Self-Adaptiveness.
While we rightfully advertise our method’s ability to dynamically adjust core decoding hyperparameters ($k$, $\alpha$, and $q$) without manual tuning, GUARD still relies on $\lambda$ and $w$. Nevertheless, our experiments suggest that neither of them substantially influences performance, and can be kept at their default values. One point to note for practical use is that we only validated this for generations of length up to 256 tokens.
6 Conclusion
We introduce GUARD, a self-adaptive decoding strategy that adaptively balances text coherence and diversity by responding to fluctuations in long-term and short-term uncertainty. By integrating a token frequency penalty into CS, GUARD reduces repetitive outputs and computational overhead while maintaining context fidelity. Theoretical analysis confirms that our estimation of global entropy preserves key statistical properties, such as unbiasedness and consistency. Comprehensive experiments across multiple open-source models, datasets, metrics, human and LLM evaluations demonstrate improvements in coherence, fluency, diversity, efficiency, informativeness, interestingness, and story development over established decoding methods. Overall, GUARD offers a robust approach for improving open-ended text generation. Future work will extend our method to additional text-generation tasks. We believe that investigating its integration with supervised fine-tuned models and exploring its performance in multilingual and low-resource settings might yield valuable insights.
Limitations
While our proposed method demonstrates improvements in open-ended text generation quality, several limitations warrant acknowledgment:
1.
Our evaluation focuses exclusively on open-ended text generation tasks. The transferability of our approach to other NLP applications, such as summarization, machine translation, and classification, remains to be investigated.
2.
The empirical evaluation is confined to English-language datasets, leaving questions about cross-lingual generalizability, particularly for low-resource languages, unanswered.
3.
While we demonstrate effectiveness across various open-source models, we have not evaluated our method on proprietary language models, which may exhibit different behaviors.
4.
Our experiments are limited to base models. The impact of our approach on supervised fine-tuned models represents an important direction for future research.
5.
Our setting was limited by a maximum length of 256 tokens, and the behaviour in long-text generation scenarios is yet to be explored.
Ethics Statement
We affirm that our research adheres to the ACL Ethics Policy. This work involves the use of publicly available datasets and does not include any personally identifiable information. For our human evaluation, we employed third-party evaluators, ensuring a rate of over \$20 per hour. An ethical concern worth mentioning is the use of language models for text generation, which may produce harmful content, either through intentional misuse by users or unintentionally due to the training data or algorithms. We declare that there are no conflicts of interest that could potentially influence the outcomes, interpretations, or conclusions of this research. All funding sources supporting this study are acknowledged in the acknowledgments section. We have diligently documented our methodology, experiments, and results, and we commit to sharing our code, data, and other relevant resources to enhance reproducibility and further advancements in the field.
Acknowledgments
This work was partially supported by the MOE Liberal Arts and Social Sciences Foundation (No.23YJAZH210), Major Program of National Social Science Foundation (No.23&ZD309), Henan Provincial Center for Outstanding Overseas Scientists (No.GZS2025004), High Level Talent International Training Program of Henan Province (No.GCC2025010), Henan Provincial Science and Technology Project (No.252102210150), Key Scientific Research Project for Universities in Henan Province (No.25A520014), and the Chinese Scholarship Council (Grant No.202308410339). Moreover, Matthias Aßenmacher received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI, under Grant No.460037581. Julian Rodemann acknowledges the funding support from the Federal Statistical Office of Germany within the co-operation project "Machine Learning in Official Statistics", the Bavarian Institute for Digital Transformation (bidt) and the Bavarian Academy of Sciences (BAdW) within a graduate scholarship. We also thank the support from the Munich Center for Machine Learning (MCML), and the Department of Statistics, LMU Munich. Last but not least, we thank all the anonymous reviewers and area chair(s) for their constructive suggestions and help throughout the review process.
Appendix
Appendix A Proofs
For ease of exposition, we restate all claims before proving them.
A.1 Proof of Proposition 1
Proposition (Unbiasedness of $H_{\text{glob}}$) Assume that the process $\{H(X)_{t}\}_{t\geq 1}$ is stationary. Then,
$\mathbb{E}\bigl{[}H_{\text{glob}}\bigr{]}=\mathbb{E}[H(X)_{t}]$
for any $\lambda\in[0,1)$.
Proof 1
Define the normalized weights
$w_{i}=\frac{\lambda^{(t-i)}}{\sum_{s=1}^{t}\lambda^{(t-s)}},\quad\text{such that}\quad\sum_{i=1}^{t}w_{i}=1,$
for any $\lambda\in[0,1)$. Then we can write the estimator (1) as
$H_{\text{glob}}=\sum_{i=1}^{t}w_{i}\,H(X)_{i}.$
Taking expectation and using linearity together with the stationarity assumption, we obtain
$\mathbb{E}\bigl{[}H_{\text{glob}}\bigr{]}=\sum_{i=1}^{t}w_{i}\,\mathbb{E}\bigl{[}H(X)_{i}\bigr{]}=\sum_{i=1}^{t}w_{i}\,H_{0}=H_{0}.$
Since under stationarity, the instantaneous entropy’s expectation at $t$ is $\mathbb{E}[H(X)_{t}]=H_{0},$ the claim immediately follows:
$E\bigl{[}H_{\text{glob}}\bigr{]}=\mathbb{E}[H(X)_{t}].$
$\square$
A.2 Proof of Proposition 2
Proof 2
We show that the variance of $H_{\text{glob}}$ vanishes as $t\to\infty$, so that by Chebyshev’s inequality the estimator converges in probability to $\mathbb{E}[H(X)_{t}]$ under the following assumptions
  1. (1)
    Stationarity and Ergodicity: The sequence $\{H(X)_{t}\}_{t\geq 1}$ is stationary and ergodic with
    $\forall t\in\{1,\dots,T\}:\;E\Bigl{[}H(X)_{t}\Bigr{]}=:H_{0}.$
  2. (2)
    Boundedness: There exists a constant $M<\infty$ such that
    $|H(X)_{t}|\leq M\quad\text{for all }t.$
  3. (3)
    Decaying Covariances: The covariances $\operatorname{Cov}\bigl{(}H(X)_{t},H(X)_{s}\bigr{)}$ decay sufficiently fast as |ts||t-s| increases.
Decaying covariances can be ensured by, e.g., strong mixing of the sequence $\{H(X)_{t}\}$ with mixing coefficients $\alpha(n)$ satisfying
$\sum_{n=1}^{\infty}\alpha(n)^{\delta/(2+\delta)}<\infty$
for some $\delta>0$.
The argument is as follows. Define the normalized weights
$w_{i}=\frac{\lambda^{t-i}}{\sum_{s=1}^{t}\lambda^{t-s}},\quad\text{such that}\quad\sum_{i=1}^{t}w_{i}=1,$
for any $\lambda\in[0,1)$.
Since
$H_{\text{glob},t}=\sum_{i=1}^{t}w_{i}\,H(X)_{t},$
its variance is given by
$\operatorname{Var}\bigl{(}H_{\text{glob},t}\bigr{)}=\sum_{i=1}^{t}\sum_{s=1}^{t}w_{i}\,w_{s}\,\operatorname{Cov}\Bigl{(}H(X)_{i},H(X)_{s}\Bigr{)}.$
Per assumptions, we have that all $H(X)_{t}$ are bounded by $M$ and their covariances decay sufficiently fast as |ts||t-s| increases. Moreover, the weights form a convex combination. Under standard mixing conditions (see, e.g., Bingham (1973)), one can show that
$\operatorname{Var}\bigl{(}H_{\text{glob},t}\bigr{)}=\mathcal{O}\Bigl{(}\frac{1}{t}\Bigr{)},$
[…]
$\mathbb{P}\Bigl{(}\bigl{|}H_{\text{glob}}-\mathbb{E}[H_{\text{glob}}]\bigr{|}>\epsilon\Bigr{)}\leq\frac{\operatorname{Var}\bigl{(}H_{\text{glob}}\bigr{)}}{\epsilon^{2}}\to 0$
where $\text{as }t\to\infty$.
Together with Proposition 1, this implies
$H_{\text{glob},t}\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\mathbb{E}[H(X)_{t}]$
as $t\rightarrow\infty$. Thus, $H_{\text{glob},t}$ is a consistent estimator.
Appendix B Further (Design Choice) Experiments
Optimal $\lambda$ Selection. We further investigate how different values of $\lambda$ influence the quality of text generations. We find that model performance remains stable for $\lambda\in\{0.91,0.93,0.95,0.97,0.99\}$, thus we select $\lambda=0.95$ as our default parameter. This experiment is based on the GPT-2 XL model (Radford et al., 2019) to balance computational expenses.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
$\lambda=0.91$ 95.25 83.64 -2.60 95.86 89.39 -2.34 96.41 83.43 -2.61
$\lambda=0.93$ 95.22 81.39 -2.60 96.31 92.79 -2.34 96.39 81.55 -2.61
$\lambda=0.95$ 95.64 85.55 -2.62 96.19 90.61 -2.35 96.35 85.83 -2.62
$\lambda=0.97$ 96.00 79.27 -2.63 96.28 92.32 -2.36 96.60 82.79 -2.63
$\lambda=0.99$ 95.95 80.48 -2.67 96.46 90.35 -2.39 96.75 79.73 -2.65
Table 6: Evaluation of performance across three datasets for $\lambda$ values in (0.91, 0.99).
Median Aggregation and Arctanh. To validate our design choices of Median aggregation and Arctanh mapping, we conducted statistical significance testing across multiple aggregation and mapping combinations. Bootstrap analysis with 95% confidence intervals demonstrated that our chosen Median + Arctanh method achieves significantly superior MAUVE scores (93.60 [92.71, 94.48]) compared to all alternative approaches ($p<0.001,t=4.98$). The effect sizes for these MAUVE improvements are consistently large (all Cohen’s $d>1.75$), with percentage improvements ranging from 3.61% to 6.39% over alternatives. While other combinations achieve better Diversity scores and certain alternatives show marginal improvements in Coherence, we follow common practices to guide these choices (Hewitt et al., 2022; Zhu et al., 2024; Nguyen et al., 2025b; Yu et al., 2024; Meister et al., 2023b). The Friedman test across all metrics ($\chi^{2}=9.10$) indicates no statistically significant differences in overall performance across methods, suggesting that trade-offs exist between optimization targets.
Aggregation Mapping Diversity MAUVE Coherence Notes
Winsorized mean Logarithmic map 95.22 88.44 -2.42 -
Median Logarithmic map 95.71 90.16 -2.48 -
Winsorized mean Arctanh 94.49 87.98 -2.35 -
Mean Logarithmic map 95.75 88.91 -2.35 -
Mean Arctanh 97.15 90.34 -2.61 -
Median Arctanh 95.20 93.60 -2.38 GUARD (Ours)
Table 7: Evaluation of different value assignment strategies and function mappings using the Wikinews dataset.
Optimal Locality Window Selection. We evaluated the performance of various locality window sizes, $w\in\{2,3,4,5,6,7,8,9\}$, with the results summarized in Table 8. The analysis indicates that a window size of $w=7$ yields good performance, while other choices in that range also display similar scores. To further validate this finding, we conducted additional experiments using multiple models, including Qwen2.5-7B, Llama2, Deepseek-llm-7B-base, Llama-3, Mistral-v0.3, Gemma-7B, and GPT2-xl. Detailed results for these experiments are provided in Tables 8, 9, 10, 11, 12, 13, and 14.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. (%)
GUARD (w=2) 93.56 87.89 -2.57 94.64 94.20 -2.35 95.94 91.25 -2.52
GUARD (w=3) 94.06 88.87 -2.67 96.17 94.29 -2.41 96.89 90.43 -2.61
GUARD (w=4) 93.11 89.32 -2.57 94.73 94.83 -2.32 95.55 93.58 -2.46
GUARD (w=5) 92.97 87.26 -2.61 96.05 93.56 -2.39 96.51 92.30 -2.54
GUARD (w=6) 92.92 86.85 -2.57 96.05 93.56 -2.34 95.69 93.57 -2.45
GUARD (w=7) 92.86 90.82 -2.61 95.20 93.60 -2.38 96.18 92.59 -2.52
GUARD (w=8) 92.51 88.29 -2.57 94.45 93.66 -2.32 95.48 91.56 -2.46
GUARD (w=9) 92.82 89.78 -2.61 94.94 94.53 -2.38 96.14 89.96 -2.52
Table 8: Qwen2.5-7B: The performance under different datasets and $w$ values shows that $w$ = 7 is an appropriate locality window. As the value of $w$ increases, the data value decreases; hence, no further experiments are conducted.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
GUARD (w=2) 95.49 90.59 -2.67 96.19 93.21 -2.25 96.21 93.08 -2.62
GUARD (w=3) 96.73 91.46 -2.87 97.18 95.33 -2.36 97.10 92.76 -2.61
GUARD (w=4) 94.73 91.27 -2.64 96.20 92.90 -2.26 96.89 91.97 -2.64
GUARD (w=5) 95.82 87.91 -2.77 96.91 92.85 -2.33 95.59 92.46 -2.59
GUARD (w=6) 94.50 90.00 -2.66 96.45 93.41 -2.24 96.15 92.78 -2.71
GUARD (w=7) 95.60 89.52 -2.75 96.76 93.34 -2.31 96.21 93.08 -2.62
GUARD (w=8) 94.74 87.78 -2.66 96.48 90.41 -2.25 96.10 92.78 -2.49
GUARD (w=9) 95.38 90.78 -2.74 96.54 94.12 -2.31 95.98 92.12 -2.59
Table 9: Llama-2: The performance under different datasets and $w$ values shows that $w$ = 7 is the appropriate locality window.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
GUARD (w=2) 95.62 89.45 -2.65 96.56 96.50 -2.21 96.32 90.36 -2.58
GUARD (w=3) 97.21 88.12 -2.74 97.55 93.36 -2.33 97.34 91.93 -2.68
GUARD (w=4) 95.41 84.49 -2.54 96.64 94.31 -2.20 96.09 90.43 -2.50
GUARD (w=5) 96.82 89.74 -2.68 97.24 94.32 -2.29 97.02 89.17 -2.61
GUARD (w=6) 95.64 83.35 -2.48 96.59 93.77 -2.20 96.34 89.46 -2.52
GUARD (w=7) 96.55 88.96 -2.65 97.14 94.84 -2.23 96.86 90.07 -2.61
GUARD (w=8) 95.70 91.09 -2.56 96.71 94.27 -2.21 96.23 88.95 -2.52
GUARD (w=9) 96.52 85.69 -2.65 97.08 94.24 -2.27 96.90 91.80 -2.61
Table 10: Deepseek-llm-7B-base: The performance under different datasets and $w$ values shows that $w$ = 7 is the appropriate locality window.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
GUARD (w=2) 95.47 89.85 -2.57 96.16 94.88 -2.15 96.02 91.20 -2.49
GUARD (w=3) 96.65 89.74 -2.66 97.25 95.49 -2.24 96.84 91.52 -2.57
GUARD (w=4) 94.94 89.98 -2.45 95.82 96.12 -2.12 95.75 90.13 -2.41
GUARD (w=5) 95.95 88.72 -2.56 96.67 93.44 -2.21 96.39 92.50 -2.52
GUARD (w=6) 94.88 90.74 -2.47 95.58 94.51 -2.14 95.73 91.14 -2.44
GUARD (w=7) 95.84 91.25 -2.55 96.45 94.94 -2.20 96.21 92.09 -2.50
GUARD (w=8) 94.77 90.92 -2.47 95.95 93.90 -2.15 95.82 91.11 -2.43
GUARD (w=9) 95.62 89.44 -2.55 96.59 92.31 -2.20 96.04 91.50 -2.48
Table 11: Llama-3: The performance under different datasets and $w$ values shows that $w$ = 7 is the appropriate locality window.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
GUARD (w=2) 94.87 92.52 -2.58 96.68 91.68 -2.20 96.25 90.89 -2.60
GUARD (w=3) 96.57 93.39 -2.71 97.49 91.89 -2.31 97.29 91.02 -2.70
GUARD (w=4) 95.10 91.64 -2.53 96.73 92.45 -2.19 96.10 91.26 -2.54
GUARD (w=5) 95.70 91.43 -2.66 97.32 89.66 -2.29 96.82 91.52 -2.65
GUARD (w=6) 94.97 91.14 -2.54 96.61 91.50 -2.21 96.20 91.57 -2.55
GUARD (w=7) 95.69 90.58 -2.63 96.85 90.49 -2.27 96.78 94.56 -2.64
GUARD (w=8) 95.05 89.26 -2.21 96.56 92.41 -2.21 96.20 92.64 -2.57
GUARD (w=9) 95.57 92.42 -2.63 97.07 93.88 -2.26 96.65 89.73 -2.62
Table 12: Mistral-v0.3: The performance under different datasets and $w$ values shows that $w$ = 7 is the appropriate locality window.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
GUARD (w=2) 95.32 89.45 -2.65 96.71 87.45 -2.16 96.28 89.39 -2.35
GUARD (w=3) 96.81 90.16 -2.62 97.53 91.44 -2.27 95.46 90.41 -2.43
GUARD (w=4) 95.86 88.22 -2.47 96.60 89.90 -2.14 96.88 89.43 -2.56
GUARD (w=5) 96.65 88.47 -2.56 97.33 89.22 -2.22 95.26 91.65 -2.67
GUARD (w=6) 95.46 90.56 -2.46 96.70 92.74 -2.15 96.44 90.16 -2.52
GUARD (w=7) 96.54 88.16 -2.55 97.23 90.76 -2.22 96.97 91.29 -2.65
GUARD (w=8) 95.59 87.81 -2.49 96.71 91.62 -2.15 96.30 91.09 -2.35
GUARD (w=9) 96.20 89.79 -2.54 97.09 88.50 -2.22 95.69 90.98 -2.64
Table 13: Gemma-7B: The performance under different datasets and $w$ values shows that $w$ = 7 is the appropriate locality window.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
GUARD (w=2) 96.64 84.21 -2.79 96.79 91.26 -2.48 97.29 84.71 -2.76
GUARD (w=3) 96.64 84.21 -2.79 96.79 91.26 -2.48 97.29 84.71 -2.76
GUARD (w=4) 94.39 83.56 -2.49 95.86 87.44 -2.25 95.56 83.13 -2.51
GUARD (w=5) 96.28 83.06 -2.66 96.49 89.83 -2.36 96.66 83.37 -2.64
GUARD (w=6) 94.60 79.72 -2.50 95.92 91.78 -2.26 95.69 81.43 -2.52
GUARD (w=7) 95.64 85.55 -2.62 96.19 90.61 -2.35 96.35 85.83 -2.62
GUARD (w=8) 94.69 84.25 -2.53 95.50 90.60 -2.55 95.68 82.50 -2.53
GUARD (w=9) 95.56 78.64 -2.62 96.17 89.35 -2.35 96.24 80.30 -2.61
Table 14: GPT2-xl: The performance under different datasets and $w$ values shows that $w$ = 7 is the appropriate locality window.
Appendix C Pseudo Code Design
Algorithm 1 GUARD Algorithm
1:
2:Prompt, Model, MaxTokens, $w$ (window size), $\lambda$ (decay factor)
3:
4:Generated text continuation
5:function GENERATETEXT(Prompt, Model, MaxTokens, w, $\lambda$)
6:  $Output\leftarrow\text{Prompt}$
7:  $entropy\_history\leftarrow[\,]$;  $tokencounts\leftarrow\{\}$;  $\varepsilon\leftarrow 1e-6$
8:  for $t=1\text{ to }MaxTokens$ do
9:   $distribution\leftarrow\text{Model.getDistribution(Output)}$
10:   $V\leftarrow\text{distribution.size}$
11:   $H_{loc}\leftarrow-\sum(\text{distribution}*\log(\text{distribution}))$ // Eq. 2
12:   $entropy\_history[t]\leftarrow H_{loc}$
13:   if $t<w$ then
[…]
15:     $\delta_{loc}\leftarrow 0$
16:   else
17:     $T\leftarrow\text{len}(entropy\_history)$
18:     $numerator\leftarrow\sum_{i=1}^{T}(\lambda^{T-i}*entropy\_history[i])$
19:     $denominator\leftarrow\sum_{i=1}^{T}(\lambda^{T-i})$
20:     $H_{glob,T}\leftarrow numerator/denominator$ //Eq. 1
21:     $med\_H\_recent\leftarrow\text{median}(entropy\_history[T-w:T])$
22:     $med\_diff\leftarrow med\_H\_recent-\text{med}(H_{glob,T})$
23:     $r_{change}\leftarrow\lvert H_{loc}-entropy\_history[T-1]\rvert/(entropy\_history[T-1]+\varepsilon)$
24:     $r_{diff}\leftarrow\lvert med\_diff\rvert/(\text{med}(H_{glob,T})+\varepsilon)$
25:     $q\leftarrow 1.0+r_{change}+r_{diff}$ //Appendix B
[…]
28:   end if
29:   $\lambda_{k}\leftarrow\lvert\delta_{loc}\rvert/(\lvert\delta_{loc}\rvert+\lvert\delta_{global}\rvert+\varepsilon)$ //Eq. 10
30:   $k_{signal}\leftarrow\exp(\lambda_{k}*\delta_{loc}+(1-\lambda_{k})*\delta_{global})$
31:   $k_{t}\leftarrow 10*(1-1/(k_{signal}+1))+5$ //Eq. 7
[…]
33:   $\alpha_{t}\leftarrow 1-1/(k_{signal\_\alpha}+1)$ //Eq. 12
[…]
35:   $scores\leftarrow\{\}$
36:   for each token $v$ in candidates do
37:     $count\leftarrow\text{tokencounts.get}(v,0)$
38:     $scores[v]\leftarrow\text{distribution}[v]*(\alpha_{t}^{\text{count}})$ //Eq. 11
39:   end for
40:   $next\_token\leftarrow\text{argmax}(scores)$
41:   $tokencounts[next\_token]\leftarrow\text{tokencounts.get}(next\_token,0)+1$
42:   $Output\leftarrow Output+next\_token$
43:  end for
44:  return $Output[\text{len(Prompt):}]$
45:end function
Appendix D Metrics
Diversity
This metric aggregates $\mathrm{n}$-gram repetition rates:
[…]
A low diversity score suggests the model suffers from repetition, and a high diversity score means the model-generated text is lexically diverse.
MAUVE
MAUVE Pillutla at al. (2021) is a metric designed to quantify how closely a model distribution $Q$ matches a target distribution $P$ of human texts. Two main types of error contribute to any discrepancy between $Q$ and $P$:
  • Type I Error: $Q$ assigns high probability to text that is unlikely under $P$.
  • Type II Error: $Q$ fails to generate text that is plausible under $P$.
These errors can be formalized using the Kullback–Leibler (KL) divergences $\mathrm{KL}(Q\!\parallel\!P)$ and $\mathrm{KL}(P\!\parallel\!Q)$. If $P$ and $Q$ do not share the same support, at least one of these KL divergences will be infinite. To address this issue, Pillutla at al. (2021) propose measuring errors through a mixture distribution
$R_{\lambda}\;=\;\lambda P\;+\;(1-\lambda)\,Q\quad\text{with}\;\lambda\in(0,1).$
This leads to redefined Type I and Type II errors given by
$\mathrm{KL}\!\bigl{(}Q\!\parallel\!R_{\lambda}\bigr{)}\quad\text{and}\quad\mathrm{KL}\!\bigl{(}P\!\parallel\!R_{\lambda}\bigr{)},$
respectively.
By varying $\lambda$ and computing these two errors, one obtains a divergence curve
$\mathcal{C}(P,Q)\;=\;\left\{\left(\exp\!\bigl{(}-c\,\mathrm{KL}(Q\!\parallel\!R_{\lambda})\bigr{)},\;\exp\!\bigl{(}-c\,\mathrm{KL}(P\!\parallel\!R_{\lambda})\bigr{)}\right)\,:\,R_{\lambda}=\lambda P+(1-\lambda)Q,\;\lambda\in(0,1)\right\},$
where $c>0$ is a hyperparameter that controls the scaling.
Finally, $\operatorname{MAUVE}(P,Q)$ is defined as the area under the divergence curve $\mathcal{C}(P,Q)$. Its value lies between 0 and 100, with higher values indicating that $Q$ is more similar to $P$.
Coherence
Proposed by Su et al. (2022), the coherence metric is defined as the averaged log-likelihood of the generated text conditioned on the prompt as:
$\operatorname{Coherence}(\hat{\boldsymbol{x}},\boldsymbol{x})=\frac{1}{|\hat{\boldsymbol{x}}|}\sum_{i=1}^{|\hat{\boldsymbol{x}}|}\log p_{\mathcal{M}}\left(\hat{\boldsymbol{x}}_{i}\mid\left[\boldsymbol{x}:\hat{\boldsymbol{x}}_{<i}\right]\right)$
where $\boldsymbol{x}$ and $\hat{\boldsymbol{x}}$ are the prompt and the generated text, respectively; [:] is the concatenation operation and $\mathcal{M}$ is the OPT model (2.7B) Zhang et al. (2022).
BERTScore
BERTScore has shown positive correlations with human judgements Zhang et al. (2020), therefore, we propose using it as an additional metric for evaluating the quality of the generated texts by each generation method with respect to the corresponding reference texts provided by human experts. By using contextual embeddings, BERTScore Zhang et al. (2020) computes a similarity score for each token in the candidate sentence with each token in the reference sentence.
Appendix E Human Evaluation
To reflect the share of human raters favoring each decoding strategy, as depicted in Table 15 in the main paper, we apply the following scoring approach:
[…]
 $\text{ Score }_{\mathrm{GUARD}}=100-\text{ Score }_{\mathrm{ACS}}.$
Dataset Coherence Fluency
ACS is better ACS and GUARD are similar GUARD is better ACS is better ACS and GUARD are similar GUARD is better
Wikitext 19% 37% 44 % 11% 67% 23%
Wikinews 24% 33% 44% 12% 67% 22%
BookCorpus 36% 27% 38% 18% 69% 13%
All 26% 32% 42% 14% 67% 19%
Table 15: Human evaluation results for ACS vs. GUARD across different datasets. Text generations are rated based on their semantic coherence and fluency.
Further, we measure the inter-rater agreement by computing Fleiss’ kappa:
$\kappa_{Fleiss}=0.41$
which reflects a moderate agreement across the human evaluators.
Figure 3: Human evaluation form, including general instructions and definitions for the evaluation criteria.
Appendix F Prompt design for LLM-as-a-judge
Instructions You are a language evaluation specialist tasked with conducting pairwise comparisons of machine-generated texts. You will compare outputs from two different decoding methods (A and B) that were generated using the same input prompt.
Evaluation Process
For each text pair, assess which version demonstrates superior quality according to six specific metrics. Your evaluation should result in one of three judgments for each metric: Method A is better Both methods are similar Method B is better
Important Note
The underlying decoding methods are randomized in their labeling as “Method A” or “Method B” across different evaluations. Do not attempt to identify patterns based on these labels, as they are arbitrarily assigned for each comparison.
Evaluation Metrics
1. Fluency
Measures how natural, smooth, and grammatically correct the text reads. Evaluates whether the language flows naturally without awkward phrasing, grammatical errors, or unnatural constructions.
2. Coherence
Assesses logical connection between ideas and overall text organization. Evaluates whether the text maintains consistent themes, follows logical progression, and avoids contradictions or non-sequiturs.
3. Factuality
Measures accuracy and truthfulness of factual claims. Evaluates whether information presented is correct and free from errors, fabrications, or misrepresentations.
4. Informativeness
Assesses the substantive content and value of the information provided. Evaluates whether the text delivers meaningful, relevant content rather than being vague, repetitive, or content-poor.
5. Interestingness
Measures how engaging, compelling, or captivating the text is. Evaluates whether the content holds attention through creativity, unique insights, or engaging stylistic elements.
6. Story Development
Assesses how effectively the narrative unfolds and progresses (where applicable). Evaluates character development, plot progression, pacing, and overall narrative structure in story-based texts.
Appendix G Further Comparative Experiments
Different Models
We examine the impact of different models on the quality of generated text, such as Llama-2, Gemma-7B, Llama-3.1, Mistral-v0.3, and Deepseek-llm-7B-base across three datasets: Wikinews, Wikitext, and BookCorpus. The results are detailed in Table 16 (diversity, MAUVE, and coherence) and Table 17 (BERTScore) reveal notable differences between the models.
Dataset Model Diversity MAUVE Coherence
CS ACS GUARD CS ACS GUARD CS ACS GUARD
Wikitext Gemma-7B 68.99 47.99 96.54 85.05 68.45 88.16 -1.50 -1.33 -2.55
Llama-3.1 9.0 81.17 95.84 27.05 75.97 91.25 -0.74 -1.52 -2.55
Mistral-v0.3 97.26 89.82 96.59 85.19 81.59 90.58 -2.09 -1.56 -2.63
Deepseek-base-7B 77.43 32.14 96.55 56.26 46.92 88.96 -2.49 -1.23 -2.66
Wikinews Gemma-7B 83.73 72.81 97.23 87.26 86.34 90.76 -1.52 -1.29 -2.22
Llama-3.1 46.84 89.66 96.45 82.36 69.44 94.94 -1.25 -1.66 -2.20
Mistral-v0.3 98.25 93.42 96.85 83.89 80.75 90.49 -1.92 -1.63 -2.27
Deepseek-base-7B 94.13 61.96 97.14 61.34 73.39 94.84 -3.69 -1.66 -2.23
BookCorpus Gemma-7B 65.65 39.76 96.97 86.65 70.59 91.29 -1.55 -1.13 -2.65
Llama-3.1 16.45 84.86 96.21 52.94 78.09 92.09 -1.00 -1.64 -2.50
Mistral-v0.3 97.27 90.74 96.78 79.03 75.50 94.56 -2.29 -1.69 -2.64
Deepseek-base-7B 62.33 42.61 96.86 62.96 47.28 90.07 -2.41 -1.24 -2.61
Table 16: Comparing GUARD with CS ($k=10,\alpha=0.6$) and ACS ($q=1$) across datasets and models of varying size. In Table 17, we also report the results of the influence of model architectures. Best results in bold.
In Table 17, we also investigate the influence of model architectures on the performance of representative contrastive-search-based decoding methods.
Dataset Model CS ACS GUARD
BERTScore (%) BERTScore (%) BERTScore (%)
Wikitext Llama-2 81.13 80.59 81.50
Gemma-7B 81.76 81.61 81.73
Llama-3.1 80.89 80.98 81.76
Mistral-v0.3 81.14 81.25 81.84
Deepseek-base-7B 80.00 80.02 81.30
Wikinews Llama-2 82.00 82.23 83.39
Gemma-7B 83.54 83.72 83.34
Llama-3.1 83.89 82.25 83.76
Mistral-v0.3 82.76 83.08 83.52
Deepseek-base-7B 77.28 80.03 83.32
BookCorpus Llama-2 79.74 81.36 81.92
Gemma-7B 81.17 80.82 81.61
Llama-3.1 81.12 81.07 81.93
Mistral-v0.3 81.56 81.43 81.83
Deepseek-base-7B 79.87 80.04 81.72
Table 17: Comparison of CS ($k=10,\alpha=0.6$), ACS ($q=1$), and GUARD in terms of BERTScore across different architectures/LLMs of varying sizes. Best results in bold.
Hyperparameter Choice for Competing Decoding Strategies
We validate our choice of hyperparameters for the competing methods (Beam search, Top-kk/Top-$p$ sampling) in Tables 18, 19, and 20.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
BB=3 15.24 32.26 -0.81 25.78 58.02 -0.91 5.48 20.23 -0.59
BB=5 14.42 26.39 -0.77 24.94 54.31 -0.87 5.39 18.69 -0.56
BB=10 12.88 23.26 -0.72 19.28 40.24 -0.77 4.02 14.17 -0.49
BB=15 12.23 20.30 -0.70 16.89 34.88 -0.72 3.32 11.93 -0.46
BB=20 11.52 17.63 -0.67 15.45 34.71 -0.69 2.94 10.68 -0.44
BB=50 8.28 13.93 -0.61 9.7 22.19 -0.58 2.24 7.60 -0.45
GUARD (Ours) 92.86 90.82 -2.61 95.20 93.60 -2.38 96.18 92.59 -2.52
Table 18: Averaged automatic evaluation results for Qwen2.5-7B for beam search (with GUARD) with different beam-widths.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
kk=3 59.01 82.40 -1.33 74.21 91.46 -1.38 47.18 87.86 -1.44
kk=5 70.69 89.23 -1.5 82.74 94.65 -1.52 67.99 90.42 -1.72
kk=10 77.30 82.91 -1.68 88.18 93.40 -1.69 82.67 91.36 -2.01
kk=15 79.66 88.95 -1.77 89.83 95.30 -1.78 85.99 94.87 -2.17
kk=20 81.62 89.73 -1.84 90.17 94.11 -1.84 88.14 92.96 -2.56
kk=50 82.69 88.97 -1.98 92.13 94.52 -2.01 91.54 93.24 -2.53
GUARD (Ours) 92.86 90.82 -2.61 95.20 93.60 -2.38 96.18 92.59 -2.52
Table 19: Averaged automatic evaluation results for Qwen2.5-7B for Top-$k$ sampling (with GUARD) with different $k$.
Method Wikitext Wikinews BookCorpus
div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh. div. (%) MAUVE (%) coh.
pp=0.60 52.43 79.56 -1.43 69.02 90.38 -1.54 38.21 80.55 -1.34
pp=0.70 64.30 83.92 -1.60 77.56 93.74 -1.65 56.81 85.18 -1.58
pp=0.80 73.80 83.80 -1.77 84.98 96.08 -1.83 72.88 90.73 -1.81
pp=0.90 79.72 88.62 -1.97 89.96 96.67 -2.02 84.76 93.30 -2.12
pp=0.95 82.47 86.37 -2.11 91.47 96.09 -2.15 88.92 94.81 -2.29
GUARD (Ours) 92.86 90.82 -2.61 95.20 93.60 -2.38 96.18 92.59 -2.52
Table 20: Averaged automatic evaluation results for Qwen2.5-7B for Top-$p$ sampling (with GUARD) with different pp.
Appendix H Case Study: Generation Examples
Prompt Returning to Australia, Headlam became Deputy Chief of the Air Staff ( DCAS ) on 26 January 1965. He was appointed a Companion of the Order of
Reference the Bath ( CB ) " in recognition of distinguished service in the Borneo Territories " on 22 June. His tenure as DCAS coincided with the most significant rearmament program the Air Force had undertaken since World War II, and with manpower shortages stemming from this expansion and from Australia’s increasing involvement in the security of South East Asia.
CS St Michael and St George ( CMG ) in the 1966 New Year Honours. What is the sentiment of this tweet? To determine the sentiment of this tweet, I will analyze the content and context of the information provided.
ACS St Michael and St George ( CMG ) on 26 June 1965. Promoted to Air Commodore on 1 July 1965, Headlam relinquished his appointment as DCAS on 15 July 1965. He was appointed Air Officer Commanding RAAF Base Amberley on 15 July 1965, a position he held until his retirement on 15 July 1967.
GUARD the Bath ( CB ) in the 1965 New Year Honours. In May 1967, he moved to London as Chief Executive of the Civil Air Lines Branch, Department of Transport. Headlam became Air Commodore-in-Chief on 30 June. On 8 November, he assumed his third and last staff command when appointed Commander-in-Chief, Far East Region. His last appointment in July–August 1970 was the appointment as Chairman and Chief Executive Officer at Qantas Airways.
Table 21: Text continuations based on prompts from the Wikitext dataset. Deviations from the context are highlighted in red. The results reveal that GUARD generates contextually relevant and fluent narratives, whereas the CS method deviates from the intended context and ACS exhibits repetitive patterns. When combined with the automatic evaluation results.
Prompt Two people have died as the result of a road traffic accident involving a truck and a car in Devon, England. The emergency services were called to the scene of
Reference a collision involving a Vauxhall Corsa and a DAF lorry at 0830 GMT on Saturday. Both occupants of the Vauxhall, a man and a woman, were pronounced dead at the scene of the collision, which was on the A361 road in the town of Knowstone, near South Molton.
CS the accident, which took place on the A379 at 1.30 p.m. Police are appealing for witnesses to come forward and help with their investigation. Anyone who saw the accident or has any information is asked to contact Devon and Cornwall Police on 101, quoting log number 466 of 20/07/2019.
ACS the accident at Bovey Tracey, Torquay, at 13:30 (14:30 CEST) on Saturday. Devon and Cornwall Police said the driver of the car, a 24-year-old man, was pronounced dead at the scene.
GUARD the accident on 8 October, near Okehampton in north Devon. Both occupants of the car, which had collided with the back of the truck, have since died.
Table 22: Text continuations based on prompts from the Wikinews dataset.
Prompt My eyes moved involuntarily from the pages of the book that I was no longer reading, to the faces of the girls who were now looking over at me. My
Reference heart began to pound in my chest, and I could hear the blood rushing in my ears. I could feel the heat of the color burning on my cheeks. I was embarrassed, no I was furious.
CS heart skipped a beat as I realized that they were looking at me with a mix of curiosity and confusion.
ACS eyes caught sight of one of the girls, and I couldn’t help but stare. "Excuse me," one of the girls said, "but what are you looking at?" I looked up at her, and then back down at the book.
GUARD heart beat faster as their gazes lingered upon me and, without even realising it myself, I was smiling back, my cheeks flushed and my eyes shining with happiness.
Table 23: Text continuations based on prompts from the BookCorpus dataset.